OpenStudy (brucebaner):

What is the missing exponent? (10^6) = 10 ^-18

OpenStudy (brucebaner):

@gerryliyana

OpenStudy (astrophysics):

-24 I guess

OpenStudy (astrophysics):

10^-24

OpenStudy (astrophysics):

If that's what you're asking...

OpenStudy (astrophysics):

\[(10^6)(10^{-24}) = 10^{6-24} = 10^{-18}\]

OpenStudy (astrophysics):

You need to be more precise with your question

OpenStudy (anonymous):

(10^6) G = 10 ^-18 G = ( 10 ^-18)/(10^6) G = 10^-18 * 10^-6 G = 10^-24

OpenStudy (anonymous):

the missing exponent is 10^-24

OpenStudy (brucebaner):

okay

OpenStudy (brucebaner):

wait i a question

OpenStudy (brucebaner):

What is the value of the expression? A. 9^-48 B. 9^-2 C. 9^2 D. 9^22

OpenStudy (brucebaner):

\[9^8/9^-4*9^10\]

OpenStudy (brucebaner):

plz help me